### v1

# Polish Academy of Sciences

Abstract
The article presents an innovative approach to the visualisation of multidimensional data, using icons inspired by Chernoff faces. The approach merges classical projection techniques with the assignment of particular data dimensions to mimic features, capitalizing on the natural ability of the human brain to interpret facial expressions. The technique is implemented as a plugin to the dpVision open-source image handling platform. The plugin allows the data to be interactively explored in the form of a swarm of ”totems” whose position in hyperspace as well as facial features represent various aspects of the data. Sample visualisations, based on synthetic test data as well as the vinhoverde 15-dimensional database on Portuguese wines, confirm the usefulness of our approach to the analysis of complex data structures.
Keywords: Multidimensional data, data visualisation, Chernoff faces, avatar glyphs, interactive 3D, dimensionality reduction, PCA, human-centred visualisation, dpVision
1 Introduction
Multidimensional data sets are a very rich resource in many applications, constantly gaining importance as both computer hardware and algorithms are increasingly capable of handling them. One element which is struggling to keep up with this development is the human visual system (and human senses in general), naturally built for the 3D space, and lacking the ability to perceive, or even imagine, a space of a higher dimensionality.
There are two ways to cope with the problem:
  1. 1.
    reduce the dimensionality of the data by taking a 3D projection or cross-section (possibly a thick-layer cross-section) of the data;
  2. 2.
    increase the dimensionality of the display by bestowing the representations of the points with additional characteristics besides their coordinates. Those additional properties can include color, size, shape parameters, any quantifiable pecularities of appearance. Each data point is then represented by a more elaborate token, which in this text will be referred to as an avatar. One well-known approach is the use of Chernoff faces [1], icons resembling a human face, with data values converted into parameters such as rounded/oblong shape of the face, hair color and length, smiling/frowning curvature of lips etc. One advantage of Chernoff faces is that they tap into the natural ability of the human observer to identify and interpret human facial traits and expressions. However, this same capacity introduces challenges, as human observers may be influenced by the emotional connotations carried by various types and expressions of human or human-like faces. Perceiving an avatar as attractive/ugly or hostile/friendly adds a factor which is not necessarily related to the actual data and can seriously distort the analysis. To overcome this problem, Lardelli [2] proposed alternative glyphs in the form of trees. Trees are very familiar to humans, but seldom evoke strong emotions.
Both approaches will be used in this work.
The type of multivariate data taken into consideration will be a cloud (finite set) of points in a multidimensional hyperspace.
The multidimensional datasets chosen for testing purposes are a) an artificial 4D shape chosen for easy visual interpretation, and b) the publicly available Vinhoverde set of data on Portuguese wines [3].
2 Symbols table
$K$
number of data points;
$n$
number of dimensions in the data to be visualized;
$m<n$
the number of dimensions the avatars are capable of representing by appearance;
$h\leq n$
number of data dimensions meaningful to the human user. By this we mean parameters which the user can relate to while thinking about the data set, because they are relevant to her work or her perception of datapoints as advantageous/disadvantageous:
$h_{s}\leq h$
number of those among them that were chosen for presentation as hyperspacial dimensions;
$h_{v}=h-h_{s}$
number of those among them that were chosen for presentation by avatar appearance;
$k\geq 3$
number of display dimensions the user is prepared to navigate. In our experiments, we limited this number to 4, so as not to overwhelm the user.
Figure 1: Data dimensions vs display dimensions
3 State of the art
Many approaches have been tested to make multivariate point clouds visible to the human observer. The human senses of vision and proprioception, and, to a lesser extent, hearing, have developed to deal with the natural 3D space, and our imagination is also strongly limited by it; the ubiquitous use of 2D screens and printed material further limits the resources at our disposal. Consequently, authors have contrived many ingenious ways of projecting (in a very broad sense of the word) multidimensional realities into a 3D or 2D visual space.
The parallel coordinate plot [4] reduces the n-dimensional coordinate system to the 2D display plane by placing the coordinate axes in parallel (usually in a vertical direction) and representing each nD data point as a polyline with individual coordinate values as its nodes (Fig. 2 (b, c)). An early application of this format in medicine was known as liver profiles, representing a single nD (typically, 6D) point as a polyline or as a set of parallel vertical bars of various height (Fig. 2 (a)). When multiple nD points need to be displayed, the polylines are colored to keep them distinguishable (Fig. 2 (b)). A smoothed version of this construct is the Andrews plot (Fig. 2 (c)) [5].
A similar approach in a different (radial) coordinate system uses radial coordinates (Fig. 2 (d)). While carrying the same information as parallel coordinates, its advantage is a somewhat easier visual comparison of plots, especially in relatively low dimensionalities (up to about 10).
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 2: Parallel and radial axis formats
A trellis chart (Fig. 3) is a compact array of small 2D plots (usually all of the same type), showing relationships between two dimensions at a time, allowing the user or reader to examine them in a more comprehensive way than if they were spread over several pages or screens. A classic setup is a full set of all (${N\choose{2}}=\frac{n(n-1)}{2}$ possible pairs of dimensions, fitted into one page or screen.
Figure 3: A trellis chart
Purdom [6] provides a review of multivariate presentation, going through the four types of variables: nominal (categorical), ordinal, discrete and continuous. Her paper includes large format trellis charts, high-resolution maps using color for a third dimension, and the use of Principal Component Analysis (PCA) to choose linear combinations of variables as aggregate ”dimensions” representing the most pertinent aspect of the data.
Figure 4: 4D hyperbox
A more condensed representation was introduced by B. Alpern and L. Carter [7]. They represent the axes of an n-dimensional coordinate system as a 2D pencil of line segments emanating from a common origin and contained in an angle of less than 180 degrees. $n-1$ additional copies of each segment are used to form a mesh of ($N\choose{2}$ paralellograms, representing the coordinate planes of the hyperspace in a hyperbox (Fig. 4), displayed with the physical appearance of a mutifaceted 3D convex polyhedron in perspective.
Despite the large number of methods proposed for the visualisation of multidimensional data, most of them focus either on mathematical reduction of dimensionality (such as PCA or t-SNE), or on geometric representations like parallel coordinates, radial plots, or static glyphs. These methods typically treat all data dimensions equally, without considering how understandable each dimension is to the human observer.
In contrast, our approach introduces two important elements that, to our knowledge, are not present in previous work. First, we clearly separate the dimensions of the dataset into two groups: those that are intuitively meaningful to users (such as quality ratings or sensory features), and those that are more abstract or technical (such as chemical coefficients). The intuitive features are assigned to visual characteristics of avatars — inspired by Chernoff faces — while the abstract features are reduced or projected into 3D spatial coordinates.
Second, we combine this concept with interactive 3D visualisation, allowing users to explore a swarm of avatars from different angles and positions. While face-based visualisation is known from Chernoff’s work, it has mostly been used in two-dimensional form. We are not aware of any previous system that integrates avatar-based glyphs with interactive spatial navigation and dimensional filtering in a 3D scene.
4 Paradigm
Quite often, when multidimensional data is mentioned, we think of abstract sets of numbers to be analyzed by pure mathematics. However, there do exist data sets in which at least a part of the parameters is intuitively recognizable and can be directly interpreted in the context of a given subject area. One example is the publicly available dataset describing Portuguese wines, where find such parameters as acidulity, sugar content, or general assessment of quality. These features are readily understandable both to the expert and to the consumer. We can therefore divide multidimensional data into two categories.
The first group is those parameters that are intuitively understandable to the human observer. The values of parameters in this group refer to perceptible qualities of the objects in question, which makes them easier to interpret and compare. In the case of wines it’s the sensory and quality attributes mentioned above, to which the user can relate from his own experience. Displaying them individually as separate dimensions provides for a better analysis and understanding of the data structure.
The other group includes those parameters, which, however important they can be mathematically, have no direct reference to human perception. They can be chemical coefficients or technical aspects of the manufacturing process, whose meaning is difficult to perceive without additional context. In such cases the computer can be used to reduce their dimensionality to a smaller number of components, preserving the most essential information while reducing the burden on the visual channel.
This conceptual division between user-recognisable and abstract dimensions is central to our method. In most visualisation systems, all dimensions are treated as equal data fields, often without considering how understandable they are to the user. While this is sufficient in purely mathematical applications, it may create problems in real-world contexts, where users expect visualisations to reflect qualities they can interpret and relate to.
By assigning the more intuitive dimensions to the appearance of avatars, we use the natural strength of human perception — the ability to notice and compare visual traits. The less intuitive dimensions are projected into 3D space using classical techniques or reduced using PCA, allowing the user to explore general patterns and clusters without being distracted by unfamiliar technical details.
This dual approach makes the system flexible and adaptable to different data types and user needs. It also reflects a real-world situation, where some features (like sweetness or colour of a wine) are easier to explain and compare, while others (like pH level or sulphate concentration) may require expert interpretation. Our system helps to balance these aspects in a single coherent visualisation.
This approach merges the advantages of two tools of analysis: on one hand, the user is offered a clear and intuitive representation of the most meaningful data fields; on the other — the less intuitive ones are filtered by the computer, finally yielding a visualisation which is both information-rich and easy to interpret.
And this representation will be delivered to the user via two faculties of the human visual system: our capability of assessing spatial relationships in a 3D space, and our aptitude for perceiving facial expressions, although in a less quantitative way.
The data dimensions which were chosen for spatial display are transformed by two matrices and one scalar parameter:
  • $\mathbf{F}$
    filtering matrix which incorporates the following functions:
    1. 1.
      Selection of named parameters;
    2. 2.
      If necessary, transformation of unnamed parameters into a reduced space (PCA);
    3. 3.
      assignment of both groups to dimensions of observation space;
  • $\mathbf{V}$
    viewing matrix, describing the position and optics of the observer moving around in observation space. This matrix and the accompanying viewing algorithm determine which data points should actually be visible in the 3D viewing space, and at what position.
  • $s$
    slab thickness is the tolerance to which a data point must intercept the viewer hyperplane to be displayed. It can be set to infinity, in which case all data points will be projected onto the hyperplane.
5 implementation
The visualisation procedure takes as its input:
  1. 1.
    the number $n$ of dimensions of the cloud;
  2. 2.
    an $n*K$ data cloud (a matrix in which each of k data points is a column, and each dimension is a row);
  3. 3.
    a transformation matrix with $n$ columns and $m+k$ rows, $m$ being the number of visual parameters of an avatar.
  4. 4.
    the $(k+2)*n$ observer matrix. It is used to transform data from the $n$-dimensional data space into the k-dimensional observation space.
    Figure 5: Fragment of dialog box assigning parameters
    These transformations, projecting the data points both into the geometric visualization hyperspace and into the parameter space of the Chernoff faces, are computed from the user’s choices made in a dialog box (Fig. 5), assigning various display forms to data dimensions. Some data dimensions can be left unassigned, and they will be subjected to PCA analysis, to reduce their number to that of the available spacial dimensions.
    One of the two additional rows, the slice thickness is used to determine the distance between a given data point and the crossection hyperplane (points too far away from it are not visualized). The other extra row is needed for homogeneous representation of the data.
  5. 5.
    the viewing matrix, part of which is controlled by the user via mouse or keyboard commands, using a multidimensional compass rose to visualize the current orientation and shift. Another user-controlled parameter is the thickness of the visible slab (area between two parallel hyperplanes).
Figure 6: A point cloud intersected by a slab
Upon completion, the procedure inserts a cloud of avatars into the viewing space, i.e. the working space of dpVision [8, 9] (see Fig. 7).
Figure 7: Example of swarm of avatars with Chernoff-like faces
6 Test data
Given the difficulty of perceiving multivariate data, we need a test data set with artificially created n-dimensional points ($n>3$) whose projection onto any combination of 3 or 2 dimensions is known and recognizable on sight.
One well-known 4D solid is the tesseract [10], a four-dimensional hypercube. However, as a regular hypersolid, it’s not the best choice for our demonstrative purposes, as its orthogonal 3D projection or cross-section in any of its constituent coordinate planes would look exactly the same. We chose to disturb this symmetry by changing its aspect ratio from $1:1:1:1$ to $3:4:6:9$, a small-integer approximation of a chain of three golden ratios. It was displayed in dpVision, and a programmed procedure rotated it around five coordinate planes (in four dimensions, rotation occurs around a plane rather than a line).
Figure 8 shows the projection, into the 3D viewing space of dpVision, of this hypersolid in one of its rotated positions.
Figure 8: The anisotropic tesseract rotating in 4D, projected into a 3D viewing space.
7 Navigating the data
The four coordinate 3D spaces of the 4D hyperspace exhibit a topology homeomorphic to that of the tetrahedron. Each of them can be transformed into each of the other ones via a $90^{\circ}$ rotation (in 4D) around one of the six coordinate planes, just like each face of a tetrahedron can be transformed into any other one by a rotation (in 3D) of approximately $71^{\circ}$ around one of the six edges. The principle is shown in Figure 9, limited to four rotations (rather than the full six) to avoid clutter. The video file accompanying this paper shows our test set performing, in steps of $10^{\circ}$ at a time, the following drill:
Figure 9: Rotating the view in 4D.
  1. 1.
    Attribute numbers to coordinate spaces:
    1. (a)
      $XYZ$
    2. (b)
      $XYT$
    3. (c)
      $XZT$
    4. (d)
      $YZT$
  2. 2.
    make first coordinate space the viewing space; place two of its axes parallel to the edges of the screen.
  3. 3.
    in steps of $10^{\circ}$, rotate the cloud in 3D $90^{\circ}$ around the axis parallel to the vertical edges of the screen, then to the horizontal.
  4. 4.
    rotate the cloud in 4D, also in nine $10^{\circ}$ steps, around one of the coordinate planes, so that the next coordinate 3D space coincides with the viewing space
  5. 5.
    repeat from step 3 until all the coordinate spaces have been viewed.
This example is only intended to demonstrate the geometric aspects of our approach. A fuller illustration uses the red wines file of the Vinhoverde wine database.
Figure 10: Wines data points
In this test Figure 10), the pH and residual sugar content were mapped to the X and Y dimensions, respectively, while dimensions Z and T represented the highest-energy plane across a hyperspace spanned by citric acid, chlorides, total sulfur dioxide, and sulphates. Overall quality of the wine was represented by the smile (or lack thereof) on the Chernoff face, and alcohol content by the length of its nose. The density (specific weight) of the wines was represented by the color of the face, and it can be observed that three distinct densities prevail in the data set, leading to a small set of colors out of a potentially near-continuous scale of hues.
8 Conclusion and Contribution
This paper introduces a new method for visualising high-dimensional data. It uses two natural abilities of the human visual system: understanding of 3D space and recognition of facial features. The method combines a 4D point cloud (rotating in 4D and projected onto a 3D space so as to make its structure accessible to human senses) with expressive icons inspired by Chernoff faces. In this way, it helps people to understand many-dimensional data in a more intuitive and user-friendly way. The 4D cloud can readily be expanded to higher dimensions, but we decided it might overwhelm the user.
The main idea is to divide the data dimensions into two groups. The first group contains features that are easy to understand — such as sensory qualities or user ratings — and these are shown as visual traits of the avatars. The second group includes more abstract or technical parameters, which are shown using 3D positions. If needed, methods such as Principal Component Analysis (PCA) are used to reduce the number of these less intuitive dimensions.
The method has been implemented as a plugin for the open-source platform dpVision. This plugin lets users explore a cloud of avatars in 3D space, where each avatar represents one data point. The user can move around the scene, change the direction of viewing, and adjust the thickness of the visible ”slab” of data, i.e. the extent, in the fourth dimension, of data that will be projected into the 3D viewing space. The assignment of data dimensions to position or appearance can also be changed.
The contribution of this work is threefold. First, it offers a new way to show complex data using both spatial position and visual appearance. Second, it provides an open and working software tool that can be used in practice. Third, it shows the method in action with two datasets: a synthetic cuboid-shaped 4D point cloud and a real-world wine quality dataset with 15 dimensions.
Our experiments show that this approach can help users to better understand complicated data structures. It can be especially useful when not all dimensions are equally important or easy to explain. Since the system is flexible and modular, it can be expanded in the future — for example, by using different avatar types or automatic selection of features.
This method can be helpful in many fields where people need to interpret data visually — such as medicine, quality control, or social sciences — especially when they are not experts in mathematics but still need to make informed decisions.
Our work proposes a new way of thinking about the structure of multidimensional data. Instead of treating all dimensions in the same way, we divide them into two categories: those that are easy to understand and those that are more technical or abstract. This division is then reflected in the visual representation. Dimensions that are more intuitive to the user are shown directly through the appearance of avatars — their shape, size, or facial expression. The other dimensions are either projected into 3D space or reduced using mathematical tools such as PCA.
We believe this distinction is important in many practical situations, where users want to understand what the data means, not just how it is mathematically structured. Unlike traditional visualisation methods, our system provides both spatial and visual channels of information, and assigns them to different types of data in a meaningful way.
To our knowledge, this is the first system that combines avatar-based glyphs with interactive 3D exploration in a way that reflects the human ability to interpret specific features more easily than others. The method has been successfully implemented in the dpVision platform and tested on both synthetic and real-world datasets. This shows that the approach is not only theoretically interesting, but also technically feasible and practically useful.
9 Future Work
There are several directions in which this work can be further developed. One natural extension is to allow for different types of avatars. While Chernoff-like faces take advantage of the human ability to perceive expressions, they also carry emotional connotations that may bias the interpretation. Therefore, alternative glyphs — such as neutral geometric shapes or biologically inspired forms like trees — could be offered as interchangeable avatar styles, depending on the context of the data and the preferences of the user.
Another area of development is the automation of dimension assignment. At present, the mapping of data dimensions to either spatial coordinates or avatar features is performed manually. In future versions, this process could be assisted or fully automated using heuristics or machine learning methods that take into account feature importance, variance, or user-defined priorities.
A further line of improvement concerns scalability. While the current implementation is effective for datasets with several hundred or a few thousand points, visual clutter becomes an issue when the number of avatars exceeds a certain threshold. Mechanisms such as dynamic filtering, clustering, or zoomable detail views could be added to improve usability for larger datasets.
Finally, a formal evaluation of the system is planned. Although the visualisation method appears promising and has shown good results in exploratory settings, its usability and effectiveness should be validated through controlled user studies in various application domains. These studies could include tasks such as anomaly detection, group comparison, or classification support, with participants from different backgrounds and levels of expertise.
Acknowledgements
The source code for the plugin is available online under the LGPL 2.1 licence at https://github.com/iitis/N-Dim-view. The source code for the dpVision framework is also available online at https://github.com/pojdulos/dpVision.
Declaration of competing interest
The authors have no competing interests to declare that are relevant to the content of this article.
Biography Section
Leszek Luchowski (D.Sc.)
has been with the IITiS PAN (formerly ZSAK PAN) since 1983, and part of its Vision Systems Group since 1987. He led the team for a part of this time, and was responsible for a number of research projects in the theory and applications of computer vision. He sustained his PhD thesis in image segmentation in 1994, and his DSc (higher doctorate) in 2014.
Dariusz Pojda (PhD)
specialises in programming (C++, Python, Java and others) and code optimisation. He is interested in data visualisation issues and is the main developer of the dpVision software. He has participated in several research projects on the application of computer vision in fields such as cultural heritage archiving, medicine, and forensics. He has also been involved in work on supporting manufacturing processes using AR/VR systems and applying 3D graphics to visualise quantum computing.