### v2

# Polish Academy of Sciences

Abstract
The article presents an innovative approach to the visualization of multidimensional data, using icons inspired by Chernoff faces. The approach merges classical projection techniques with the assignment of particular data dimensions to mimic features, capitalizing on the natural ability of the human brain to interpret facial expressions. We introduce a semantic division of data dimensions into intuitive and technical categories, assigning the former to avatar features and projecting the latter into a hyperspace of four, or potentially more dimensions. The technique is implemented as a plugin to the dpVision open-source image handling platform. The plugin allows the data to be interactively explored in the form of a swarm of avatars whose position in hyperspace as well as facial features represent various aspects of the data. Sample visualizations, based on synthetic test data as well as the 12-dimensional database on Portuguese Vinho Verde wines, confirm the usefulness of our approach to the analysis of complex data structures.
1 Introduction
Multidimensional data sets are a very rich resource in many applications, constantly gaining importance as both computer hardware and algorithms are increasingly capable of handling them. One element which is struggling to keep up with this development is the human visual system (and human senses in general), naturally built for the 3D space, and lacking the ability to perceive, or even imagine, a space of a higher dimensionality.
There are two ways to cope with the problem:
  1. 1.
    reduce the dimensionality of the data by taking a 3D projection or cross-section (possibly a thick-layer cross-section) of the data;
  2. 2.
    increase the dimensionality of the display by bestowing the representations of the points with additional characteristics besides their coordinates. Those additional properties can include color, size, shape parameters, any quantifiable visual peculiarities. Each data point is then represented by a more elaborate token, which in this text will be referred to as an avatar. One well-known approach is the use of Chernoff faces [1], icons resembling a human face, with data values converted into parameters such as rounded/oblong shape of the face, hair color and length, smiling/frowning curvature of lips etc. One advantage of Chernoff faces is that they tap into the natural ability of the human observer to identify and interpret human facial traits and expressions. However, this capacity also introduces challenges, as human observers may be influenced by the emotional connotations carried by various types and expressions of human or human-like faces. Perceiving an avatar as attractive/ugly or hostile/friendly adds a factor which is not necessarily related to the actual data and can seriously distort the analysis. To overcome this problem, Lardelli [2] proposed alternative glyphs in the form of trees. Trees are very familiar to humans, but seldom evoke strong emotions.
In this work, we propose a novel approach that semantically separates data dimensions into two groups: those that are intuitively meaningful to the user (e.g., sensory, categorical, interpretable parameters), and those that are abstract or technical. Users may assign meaningful dimensions to either geometric space or avatar visual features. If fewer than the required number of dimensions are assigned, the system automatically supplements them using anonymous dimensions, which are processed via PCA to generate missing spatial or visual features. This approach is implemented as an interactive plugin for the dpVision platform [3, 4], allowing users to explore a cloud of 3D avatars positioned in a 4D data space and visualized through dynamic projections into 3D space (see Figure 1). By merging spatial navigation with avatar-based representation, our plugin provides a human-centered and intuitive visualization of complex multidimensional data.
Figure 1: Data dimensions vs display dimensions
To validate the proposed approach, we use two types of multivariate datasets: a) an artificial 4D structure designed for easy geometric interpretation, and b) the publicly available dataset containing 12-dimensional data on Portuguese Vinho Verde wines [5, 6].
2 STATE OF THE ART
Many approaches have been tested to make multivariate point clouds visible to the human observer. The human senses of vision and proprioception, and, to a lesser extent, hearing, have developed to deal with the natural 3D space, and our imagination is also strongly limited by it; the ubiquitous use of 2D screens and printed material further limits the resources at our disposal. Consequently, authors have contrived many ingenious ways of projecting (in a very broad sense of the word) multidimensional realities into a 3D or 2D visual space.
The parallel coordinate plot [7] reduces the n-dimensional coordinate system to the 2D display plane by placing the coordinate axes in parallel (usually in a vertical direction) and representing each nD data point as a polyline with individual coordinate values as its nodes (Figure 2 (b, c)). An early application of this format in medicine was known as liver profiles, representing a single nD (typically, 6D) point as a polyline or as a set of parallel vertical bars of various height (Figure 2 (a)). When multiple nD points need to be displayed, the polylines are colored to keep them distinguishable (Figure 2 (b)). A smoothed version of this construct is the Andrews plot (Figure 2 (c)) [8].
A similar approach in a different (radial) coordinate system uses radial coordinates (Figure 2 (d)). While carrying the same information as parallel coordinates, its advantage is a somewhat easier visual comparison of plots, especially in relatively low dimensionalities (up to about 10).
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 2: Parallel and radial axis formats
A trellis chart (Figure 3) is a compact array of small 2D plots (usually all of the same type), showing relationships between two dimensions at a time, allowing the user or reader to examine them in a more comprehensive way than if they were spread over several pages or screens. A classic setup is a full set of all $\binom{N}{2}=\frac{n(n-1)}{2}$ possible pairs of dimensions, fitted into one page or screen.
Figure 3: A trellis chart
Purdom [9] provides a review of multivariate presentation, going through the four types of variables: nominal (categorical), ordinal, discrete, and continuous. Her paper includes large format trellis charts, high-resolution maps using color for a third dimension, and the use of Principal Component Analysis (PCA) to choose linear combinations of variables as aggregate "dimensions" representing the most pertinent aspect of the data.
Figure 4: 4D hyperbox
A more condensed representation was introduced by B. Alpern and L. Carter [10]. They represent the axes of an n-dimensional coordinate system as a 2D pencil of line segments emanating from a common origin and contained in an angle of less than 180 degrees. The $n-1$ additional copies of each segment are used to form a mesh of $\binom{N}{2}$ parallelograms, representing the coordinate planes of the hyperspace in a hyperbox (Figure 4), displayed with the physical appearance of a multifaceted 3D convex polyhedron in perspective.
Despite the large number of methods proposed for the visualization of multidimensional data, most of them focus either on mathematical reduction of dimensionality (such as PCA or t-SNE), or on geometric representations like parallel coordinates, radial plots, or static glyphs. These methods typically treat all data dimensions equally, without considering how understandable each dimension is to the human observer.
An alternative, cognitively grounded approach was proposed by Pflughoeft et al. [11], who introduced data avatars — metaphorical figures designed to visually encode multidimensional data in ways that align with human intuitive perception. Their work, based on dual-process cognitive theory, shows that representations tailored to intuitive understanding can improve user performance in data interpretation tasks.
We propose an approach that introduces two complementary elements not present in previous work. First, we clearly separate the dimensions of the dataset into two groups: those that are intuitively meaningful to users (such as quality ratings or sensory features), and those that are more abstract or technical (such as chemical coefficients). The intuitive features are assigned to visual characteristics of avatars—inspired by Chernoff faces—while the abstract features are reduced or projected into 3D spatial coordinates.
Second, we combine this concept with interactive 3D visualization, allowing users to explore a swarm of avatars from different angles and positions. While face-based visualization is known from Chernoff’s work, it has mostly been used in two-dimensional form. We are not aware of any previous system that integrates avatar-based glyphs with interactive spatial navigation and dimensional filtering in a 3D scene.
3 NOTATION AND TERMINOLOGY
Throughout this paper, we use consistent notation to describe data dimensions, transformation matrices, and visualization parameters. The key symbols and abbreviations are summarized in Table 1, and are referenced throughout the description of the projection procedure and implementation details.
Table 1: Symbols and notation used throughout the paper.
Symbol Meaning Constraints
Input data parameters
$N$ Total number of data points
$n$ Total number of data dimensions
$X$ Original data matrix $X\in\mathbb{R}^{n\times N}$
Implementation-dependent mapping
$k$ Number of navigable display dimensions $k\geq 3$ (typically 4)
$m$ Number of available avatar visual features $m\leq n-k$ (currently 10)
User’s mapping
$h_{S}$ Number of human-meaningful dimensions mapped to space $h_{S}\leq k$
$h_{V}$ Number of human-meaningful dimensions mapped to avatars $h_{V}\leq m$
$h_{A}$ Number of anonymous dimensions $h_{A}\leq n-(h_{S}+h_{V})$
Projection matrices
$F$ Filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$
$V$ View matrix $V\in\mathbb{R}^{(k+1)\times(k+1)}$
Derived datasets
$X_{\text{spatial}}$ Spatial projection of $X$ $X_{\text{spatial}}\in\mathbb{R}^{k\times N}$
$X_{\text{visual}}$ Avatar feature mapping of $X$ $X_{\text{visual}}\in\mathbb{R}^{m\times N}$
$X_{\text{view}}$ Spatial data after applying $V$ $X_{\text{view}}\in\mathbb{R}^{k\times N}$
4 PARADIGM
Multidimensional data sets often contain a mixture of parameters: some of them intuitively understandable to humans, and others more abstract or technical. To maximize interpretability, we propose a semantic division of data dimensions into two groups:
  • Intuitive dimensions, such as quality ratings or sensory properties, are mapped onto visual characteristics of avatars (e.g., face shape, expressions, color).
  • Technical dimensions, such as chemical or structural parameters, are projected into spatial coordinates in a 3D observation space after optional dimensionality reduction.
Figure 5: Conceptual diagram illustrating the separation of multidimensional data into spatial coordinates and avatar visual features through filtering and viewing transformations.
Let the original data set be represented by a matrix $X\in\mathbb{R}^{n\times N}$, where $n$ is the number of original dimensions and $N$ is the number of data points.
A filtering matrix $F\in\mathbb{R}^{(k+m)\times n}$ is constructed to separate and process the dimensions, where $k$ is the number of spatial dimensions and $m$ the number of visual avatar features. The filtered data are:
$X_{\text{filtered}}=F\cdot X$
Here, each row of $F$ defines a new output feature (spatial or visual) as a linear combination of the original dimensions of $X$.
Spatial coordinates are obtained from a subset $F_{\text{spatial}}$, where $F_{\text{spatial}}\in\mathbb{R}^{k\times n}$, resulting in:
$X_{\text{spatial}}=F_{\text{spatial}}\cdot X$
Similarly, visual parameters are derived from $F_{\text{visual}}$, where $F_{\text{visual}}\in\mathbb{R}^{m\times n}$
$X_{\text{visual}}=F_{\text{visual}}\cdot X$
The spatial data are further transformed using a viewing matrix $V\in\mathbb{R}^{k\times n}$ (where typically $k=3$ or $4$):
$X_{\text{view}}=V\cdot X_{\text{spatial}}$
This dual mapping—avatars for intuitive features, spatial positioning for abstract features—provides a flexible and human-centered way to visualize high-dimensional data.
The conceptual framework of the proposed method, including the separation of intuitive and technical dimensions and the application of the filtering and viewing matrices, is illustrated in Figure 5.
5 IMPLEMENTATION
Based on the conceptual framework presented in the previous section, the practical implementation of the proposed visualization method involves the construction of appropriate transformation matrices, dimensionality reduction where necessary, and the assignment of selected data dimensions to spatial coordinates or avatar visual features.
The implementation uses the Eigen library (C++), which allows efficient operations on matrices and vectors.
The input data are stored in a matrix of size $n\times N$:
Eigen::MatrixXd X(n, N);
Each column corresponds to a single point in the data space $\mathbb{R}^{n}$, and each row represents one of the $n$ dimensions in that space.
5.1 Assigning dimensions to categories
Note: To avoid confusion, we refer to dimensions when describing input data, and to features when referring to the visualized output.
Figure 6: Dialog window for assigning input dimensions to feature categories: spatial, visual, unnamed, or skipped. For spatial and visual categories, a specific output feature label can be selected.
Input dimensions can be assigned to one of three feature categories: spatial, visual, or anonymous. Additionally, any dimension may be marked as skipped and excluded from further processing.
This assignment is performed through a dedicated dialog window (Figure 6), which enables the user to associate input dimensions with specific feature categories and, when applicable, select predefined visual or spatial feature labels.
The complete set of available labels for spatial and visual features is presented in Table 2.
The resulting configuration is stored in index tables. It is not required to assign all available features — any missing spatial or visual values are automatically interpolated during subsequent processing.
Table 2: List of features used in the implementation
Category Feature Description
User-defined spatial features
spatial X X coordinate in 4D space
spatial Y Y coordinate in 4D space
spatial Z Z coordinate in 4D space
spatial T T coordinate in 4D space
User-defined avatar facial attributes
visual Skin_C avatar’s skin color
visual Hair_C avatar’s hair color
visual Eye_S distance between the avatar’s eyes
visual Nose_L avatar’s nose length
visual Mouth_W avatar’s mouth width
visual Smile degree of curvature of the avatar’s smile
visual Frown level of frowning expression
visual Hair_L avatar’s hair length
visual Face_Elong avatar’s face proportions
visual Iris_C avatar’s iris color
features not defined by the user
anonymous used for principal component analysis (PCA)
Features selected by the user to skip
skipped excluded from further processing
5.2 Filtering matrix
The first step is to create the filtering matrix, which in the implementation has the following form:
Eigen::MatrixXd F(k + m, n);
F.setZero();
To simplify later operations, we assume that the first $k$ rows of this matrix correspond to spatial features (in our implementation, these are the 4 features: X, Y, Z, T), and the following $m$ rows correspond to visual features (in our case: 10 features representing details of the avatar’s appearance). The matrix $F$ is initially filled with zeros and then completed based on the feature assignments:
  • First, for each of the $h_{S}$ spatial features and $h_{V}$ visual features selected by the user, 1 is set at the intersection of the row for that feature and the column of the original input dimension.
  • Then, if there are still unused rows in $F$, and the user has selected anonymous features ($h_{A}>0$), a PCA is performed on the input dimensions assigned to those anonymous features. The $(k+m)-(h_{S}+h_{V})$ strongest components are selected, and their coefficient vectors are inserted into the remaining empty rows of $F$.
The transformed data is calculated as the product of the filtering matrix and the input matrix, followed by normalization. The procedure for creating the filtering matrix and transforming the input data is shown in Algorithm 1.
Algorithm 1 Implementation of Data Projection
0:  $X$ (all features), $Asgn$ (assignment of features to groups)
0:  $X_{\text{filtered}}$ (projected data)
1:  Initialize $F$ (zero matrix)
2:  for each feature index $j$ do
3:     if assignment of feature $j$ is ignored then
4:        Skip
5:     else if assignment of feature $j$ is spatial then
6:        Mark corresponding spatial row in $F$
7:     else if assignment of feature $j$ is visual then
8:        Mark corresponding visual row in $F$
9:     else if assignment of feature $j$ is anonymous then
10:        Save index of feature $j$ for PCA
11:     end if
12:  end for
13:  if number of mapped features $(h_{S}+h_{V})<(k+m)$ then
14:     Perform SVD on anonymous features
15:     Select $(k+m)-(h_{S}+h_{V})$ top components
16:     Insert principal directions into empty rows in $F$
17:  end if
18:  $X_{\text{filtered}}=F\times X$
19:  for each row in $X_{\text{filtered}}$ do
20:     Center and scale (standardize)
21:     if row corresponds to spatial dimension then
22:        Shift by $+0.5$
23:     end if
24:  end for
25:  return  $X_{\text{filtered}}$
In this way, we obtain a table of $N$ points, each with $(k+m)$ dimensions. These dimensions can be divided into two categories:
  • The first $k$ rows form the matrix X_spatial, which contains spatial features.
  • The following $m$ rows form the matrix X_visual, which contains avatar visual features.
Eigen::MatrixXd X_filtered(k + m, N);
X_filtered = F * X;
Eigen::MatrixXd X_spatial = X_filtered.topRows(k);
Eigen::MatrixXd X_visual = X_filtered.bottomRows(m);
The matrix $\mathbf{X}_{\text{spatial}}$ is now a ready-to-use $k$-dimensional dataset for further processing.
5.3 View matrix
The implementation supports navigation in a $k$-dimensional space (typically 4D) using affine transformations. These transformations are represented by a square matrix of size $(k+1)\times(k+1)$, which combines both rotation and translation.
The core of this matrix is a $k\times k$ rotation matrix, generated by specifying the rotation plane (two coordinate axes) and the rotation angle.
In the current plugin implementation, all 4D rotations are carried out in coordinate planes such as XY, XZ, XT, YZ, YT, or ZT. These correspond to classic SO(2) rotations within two-dimensional subspaces. However, the implementation can be extended to support full SO(4) rotations, offering a more expressive and flexible navigation model.
The remaining part of the view matrix stores a translation vector and a homogeneous row to enable affine transformations.
In code, the matrix can be built as follows:
// Rotation only (k x k)
Eigen::MatrixXd R = generateRotationMatrix(k, axis1, axis2, angle);
// Create full (k+1 x k+1) view matrix with translation
Eigen::MatrixXd V = Eigen::MatrixXd::Identity(k + 1, k + 1);
V.topLeftCorner(k, k) = R;
V.topRightCorner(k, 1) = translation_vector; // optional
To apply the transformation, the spatial data must be converted into homogeneous coordinates by adding a row of ones:
Eigen::MatrixXd X_hom(k + 1, N);
X_hom.topRows(k) = X_spatial;
X_hom.row(k) = Eigen::RowVectorXd::Ones(N);
Eigen::MatrixXd X_view = V * X_hom;
This allows smooth 4D navigation, including both rotation and panning of the data in the observation space.
5.4 Slab-based Filtering
To limit the number of visible data, a slab-based filtering mechanism is used. The goal is to retain only those data points that lie close to the viewing space—within a given distance slab_thickness from a hyperplane defined by the current rotation.
The filtering uses a normal vector to the viewing space to compute signed distances from each point. Depending on whether the filtering is applied before or after transforming the data by the view matrix, different normal vectors are used:
  • if filtering is done in the original spatial coordinates (X_spatial), the normal vector can be taken as the last row of the rotation matrix, which defines the orientation of the view;
  • if filtering is done after applying the rotation (X_view), the normal vector can be defined in canonical coordinates as $[0,0,0,1]$, which corresponds to the last axis.
Each point is projected onto the normal vector, and only those within the allowed threshold are retained.
Eigen::VectorXd slab_values = normal_vector.transpose() * X_view;
auto mask = slab_values.cwiseAbs().array() < slab_threshold;
Eigen::MatrixXd X_visible = skip_masked(X_view, mask);
Points that satisfy the slab condition—meaning they are sufficiently close to the viewing hyperplane—are passed to the next stage of the pipeline. This filtering effectively creates a thick slice through the multidimensional space (Figure 7).
Figure 7: A cloud of points intersected by a slab, using a threshold of 1.5. Red points that lie outside the slab are discarded.
5.5 3D projection
The four-dimensional coordinates are then projected into 3D space using a standard perspective projection. Avatars are generated as three-dimensional objects based on visual features stored in the $\mathbf{X}_{\text{visual}}$ matrix and rendered using the tools provided by the dpVision platform (Figure 8). To ensure interactive performance with large datasets, the avatars are deliberately simplified in geometry, allowing a large number of them to be rendered and manipulated in real time.
Figure 8: Part of a sample swarm of avatars with Chernoff-like faces perspectively projected into 3D space.
6 TEST DATA
Given the difficulty of perceiving multivariate data, we need a test data set with artificially created n-dimensional points ($n>3$) whose projection onto any combination of 3 or 2 dimensions is known and recognizable on sight.
One well-known 4D solid is the tesseract [12], a four-dimensional hypercube. However, as a regular hypersolid, it’s not the best choice for our demonstrative purposes, as its orthogonal 3D projection or cross-section in any of its constituent coordinate planes would look exactly the same. We chose to disturb this symmetry by changing its aspect ratio from $1:1:1:1$ to $3:4:6:9$, a small-integer approximation of a chain of three golden ratios. It was displayed in dpVision, and a programmed procedure rotated it around five coordinate planes (in four dimensions, rotation occurs around a plane rather than a line).
Figure 9 shows the projection, into the 3D viewing space of dpVision, of this hypersolid in one of its rotated positions.
Figure 9: The anisotropic tesseract rotating in 4D, projected into a 3D viewing space.
7 NAVIGATING THE DATA
The four coordinate 3D spaces of the 4D hyperspace exhibit a topology homeomorphic to that of the tetrahedron. Each of them can be transformed into each of the other ones via a $90^{\circ}$ rotation (in 4D) around one of the six coordinate planes, just like each face of a tetrahedron can be transformed into any other one by a rotation (in 3D) of approximately $71^{\circ}$ around one of the six edges. The principle is shown in Figure 10, limited to four rotations (rather than the full six) to avoid clutter. The video file accompanying this paper shows our test set performing, in steps of $10^{\circ}$ at a time, the following drill:
Figure 10: Rotating the view in 4D.
  1. 1.
    Attribute numbers to coordinate spaces:
    1. (a)
      $XYZ$
    2. (b)
      $XYT$
    3. (c)
      $XZT$
    4. (d)
      $YZT$
  2. 2.
    make first coordinate space the viewing space; place two of its axes parallel to the edges of the screen.
  3. 3.
    in steps of $10^{\circ}$, rotate the cloud in 3D $90^{\circ}$ around the axis parallel to the vertical edges of the screen, then to the horizontal.
  4. 4.
    rotate the cloud in 4D, also in nine $10^{\circ}$ steps, around one of the coordinate planes, so that the next coordinate 3D space coincides with the viewing space
  5. 5.
    repeat from step 3 until all the coordinate spaces have been viewed.
This example is only intended to demonstrate the geometric aspects of our approach. A fuller illustration uses the red wines file of the Vinho Verde wine dataset.
Figure 11: Example of visualization of the data from the Vinho Verde wine dataset.
In this test (Figure 11), the pH and residual sugar content were mapped to the X and Y dimensions, respectively, while dimensions Z and T represented the highest-energy plane across a hyperspace spanned by citric acid, chlorides, total sulfur dioxide, and sulfates. Overall quality of the wine was represented by the smile (or lack thereof) on the Chernoff face, and alcohol content by the length of its nose. The density (specific weight) of the wines was represented by the color of the face, and it can be observed that three distinct densities prevail in the data set, leading to a small set of colors out of a potentially near-continuous scale of hues.
8 CONTRIBUTIONS
This work provides several contributions to the field of multidimensional data visualization:
  1. 1.
    Semantic division of data dimensions
    We propose a novel conceptual framework that divides the dimensions of a dataset into two categories:
    • dimensions that are intuitively understandable to human users (e.g., quality ratings, sensory characteristics),
    • and more abstract or technical dimensions (e.g., chemical or process parameters).
    Unlike previous methods, which typically treat all dimensions equally, we assign intuitive dimensions to visual features of avatars (shape, facial expression, color) and project abstract dimensions into 3D spatial coordinates after possible dimensionality reduction.
  2. 2.
    Integration of avatar-based glyphs with interactive 3D navigation
    While Chernoff faces have historically been used in static 2D representations, we integrate avatar glyphs into a fully interactive 3D environment. Users can freely explore the data cloud, rotate the viewpoint, adjust the visible slab, and examine clusters and patterns dynamically.
  3. 3.
    Practical implementation as an open-source plugin
    We implemented our method as a working plugin for the open-source platform dpVision. The tool supports user-driven mapping of dimensions to visual properties and interactive spatial exploration, offering immediate practical applicability.
  4. 4.
    Demonstration on both synthetic and real-world datasets
    We validated our approach using two datasets:
    • a synthetic 4D anisotropic hypercube (for controlled geometric validation),
    • and the Vinho Verde wine dataset (a 12-dimensional real-world dataset), illustrating the practical benefits of the method in an applied context.
  5. 5.
    Foundations for future extensibility
    The visualization system is designed with modularity and flexibility in mind. Its architecture supports alternative avatar types, adjustable mappings of data dimensions, and dynamic filtering. While this version targets moderately sized datasets, its structure provides a basis for scaling to larger data collections, as further discussed in the Future Work section.
9 CONCLUSION
In this paper, we presented a novel method for visualizing high-dimensional data by combining two natural strengths of the human visual system: the understanding of 3D space and the recognition of facial traits.
Our approach introduces an important distinction between intuitive and abstract data dimensions and maps them differently for optimal interpretability: intuitive features are assigned to avatar characteristics, while abstract features are reduced and projected into 3D space.
We implemented the method as a plugin for the open-source dpVision platform and demonstrated its effectiveness on synthetic and real-world datasets. Preliminary results suggest that this combination of avatar-based glyphs and interactive 3D navigation offers an intuitive and flexible way to explore complex data structures, particularly when users are not domain experts in mathematics.
Further development directions are outlined in the next section and include formal usability studies, exploration of alternative glyph styles to mitigate emotional bias, automated feature mapping, and scalability improvements to handle larger datasets.
By distinguishing between different types of dimensions and tailoring their visual representation, our plugin proposes a human-centered paradigm that could benefit many fields where complex data must be understood and interpreted visually.
10 FUTURE WORK
There are several directions in which this work can be further developed. One natural extension is to allow for different types of avatars. While Chernoff-like faces take advantage of the human ability to perceive expressions, they also carry emotional connotations that may bias the interpretation. Therefore, alternative glyphs—such as neutral geometric shapes or biologically inspired forms like trees—could be offered as interchangeable avatar styles, depending on the context of the data and the preferences of the user. Such alternatives could reduce the emotional bias associated with human-like faces and allow more neutral data interpretation depending on the application domain.
Another area of development is the automation of dimension assignment. At present, the mapping of data dimensions to either spatial coordinates or avatar features is performed manually. In future versions, this process could be assisted or fully automated using heuristics or machine learning methods that take into account feature importance, variance, or user-defined priorities.
A further line of improvement concerns scalability. While the current implementation is effective for moderately sized datasets (typically up to a few thousand points), its perceptual clarity decreases as the number of avatars grows. Glyph-based methods such as ours are not ideally suited for very large-scale datasets, as visual clutter and cognitive overload quickly become limiting factors. Future versions could integrate mechanisms such as dynamic filtering, clustering, sampling, or zoomable detail views to maintain interpretability in larger data collections. Hybrid solutions that combine glyphs with statistical overviews may also be explored.
Finally, a formal evaluation of the system is planned. Although the visualization method appears promising and has shown good results in exploratory settings, its usability and effectiveness should be validated through controlled user studies in various application domains. These studies could include tasks such as anomaly detection, group comparison, or classification support, with participants from different backgrounds and levels of expertise.
11 SUPPLEMENTARY MATERIAL
The source code for the plugin is available online under the LGPL 2.1 licence at https://github.com/iitis/N-Dim-view. The source code for the dpVision framework is also available online at https://github.com/pojdulos/dpVision.